OpenStudy (anonymous):

Simplify 6(x + y) + (x - y).

OpenStudy (anonymous):

@AG23 not a option

OpenStudy (tommynaut):

Expand the brackets. What do you get? Then, collect like terms. I'll give you a similar example. 4(x-y) + (x-2y) = 4x - 4y + x - 2y = 4x + x - 4y - 2y = 5x - 6y Try using a similar approach for your question

OpenStudy (anonymous):

Distribute 6(x-y)

OpenStudy (anonymous):

6(x+y)*

OpenStudy (anonymous):

i dont no how to do the whole adding and subtracting letters thing without the little guide

OpenStudy (anonymous):

7x+5y

OpenStudy (anonymous):

explain?

OpenStudy (usukidoll):

please refrain from giving direct answers as it is against OpenStudy's Code of Conduct. The purpose of this site is to guide the user to answering their question.

OpenStudy (triciaal):

to expand the brackets and to distribute means the same thing. multiply the term outside by each term inside follow the example given by @ Tommynaut above

OpenStudy (tommynaut):

You can add and subtract letters (also known as pronumerals or variables) that are the same. For example, I can't do anything with x + 2y, because the letters are different! How about 2x + 3x? The letters are the same! So, we just add the COEFFICIENTS (the numbers that are next to and multiplied with the letters) together. So, for 2x + 3x, we do 2+3 = 5. So 2x + 3x = 5x. Remember to keep the x!

OpenStudy (anonymous):

So basically 6(x+y). Is the same as (6x)+(6y) Do you get it now?

OpenStudy (anonymous):

Simplifying 6(x + y) + (x + -1y) (x * 6 + y * 6) + (x + -1y) (6x + 6y) + (x + -1y) Remove parenthesis around (x + -1y) 6x + 6y + x + -1y Reorder the terms: 6x + x + 6y + -1y Combine like terms: 6x + x = 7x 7x + 6y + -1y Combine like terms: 6y + -1y = 5y 7x + 5y So like this? @UsukiDoll sorry Im new

OpenStudy (tommynaut):

It's sort of like saying, let x mean balls. I have 2 balls, and another 3 balls. That means I have 2 balls + 3 balls, which is 5 balls.

OpenStudy (anonymous):

there is no other number for me to do that with..... (ex: the whole (x-y) thing is confusing because i dont no what to do about it)

OpenStudy (anonymous):

@Tommynaut i do understand what you are doing though

OpenStudy (usukidoll):

@Frascamr it's ok... but please refer to this website for more information about the code of conduct. http://openstudy.com/code-of-conduct As for the explanation. I think it's ok, but a bit too much. try step by step :)

OpenStudy (triciaal):

when you don't see the number it is understood the number is 1

OpenStudy (anonymous):

oh

OpenStudy (tommynaut):

Imagine that there's a 1 outside (x-y), because really, 1 is multiplied by everything all the time. The number 6? Yeah, that's just 6 times 1. So (x-y) can be thought of as 1(x-y), which, if you expand the brackets, just gives you x - y

OpenStudy (anonymous):

i got 5x + 4Y which i no is wrong because it isnt a option

OpenStudy (tommynaut):

Ok, let's go through step-by-step. So we start with... 6(x + y) + (x - y). Now what? Let's expand the brackets. Do you know how to do that?

OpenStudy (anonymous):

yea you put it as 6x - 6y + 1x - 1y ....... i think

OpenStudy (usukidoll):

everything is right so far except the sign for 6(x+y)

OpenStudy (anonymous):

ops

OpenStudy (tommynaut):

Almost! But it's 6x + 6y, not 6x - 6y, where did the negative sign come from?

OpenStudy (usukidoll):

typos can happen XD

OpenStudy (anonymous):

i had wrote it on the original problem then marked it out and then did it again on the second one >.<

OpenStudy (usukidoll):

6x + 6y + x - y . now combine like terms

OpenStudy (anonymous):

oh now i see the answer well at least half 7x unless 7x is the answer

OpenStudy (usukidoll):

you can rearrange the equation to make the computation easier 6x + x + 6y - y .

OpenStudy (anonymous):

12xy?

OpenStudy (usukidoll):

nooooooooooo... we just combine like terms. we can't combine x and y together

OpenStudy (anonymous):

so just 12?

OpenStudy (anonymous):

oh i no what you meant

OpenStudy (usukidoll):

6x + x + 6y - y I'm going to factor out the variables to make it easier x(6+1)+y(6-1)

OpenStudy (anonymous):

6x + 6y

OpenStudy (anonymous):

7x + 7y

OpenStudy (usukidoll):

almost... x is correct y isn't what's 6-1

OpenStudy (anonymous):

5

OpenStudy (usukidoll):

yes so we have 5y so our answer is 7x+5y

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

thanks